OpenStudy (kj4uts):

If f(x)=4x, g(x)=5x-3, and h(x)=√x, find (fogoh)(4)? Is this how the problem is solved: (fogoh)=4(5(sqrt x)-3) =20 sqrt x-12 x=4, fogoh= 40-12= 28

OpenStudy (kj4uts):

Problem attached below! Please help and explain!

OpenStudy (zzr0ck3r):

you got it

OpenStudy (zzr0ck3r):

\((f\circ g\circ h)(x) =f\circ(g(h(x)))=f(g(h(x))) = 4(5\sqrt{x}-3) \)